OpenStudy (anonymous):

I need help understanding how to determine whether the graphs of equations are parallel,perpendicular,or neither.

OpenStudy (anonymous):

\[y=x+11 and y=-x+2\] These are teh two equations but I dont understand how to find the slope to determine which they are.

OpenStudy (anonymous):

Thank you!!!

OpenStudy (phi):

In case you did not figure it out, when you see an equation like y= x+11 that is the equation of a line in slope-intercept form y= mx+b the m is the slope. The m is the number that multiplies the x in your case, m is 1 y= 1x +11 (people don't bother to write the 1 because 1*x= x) for the other equation y= -x+2 this is also in y= mx+b form. m for this equation is -1. The slope is -1 perpendicular lines have "minus flipped" slopes. Example: if the slope of one line is \( -\frac{1}{2} \), a line perpendicular to it will have slope +2 (flip the -1/2 to get -2, then minus it - -2= +2) For your problem, slopes 1 and -1 pass perpendicular the test: 1/1 (flip 1) then minus: -1/1 which is just -1.